# Beyond Aesthetics: Cultural Competence in Text-to-Image Models

### v5

@@ 1 Introduction
"Technology is not merely a system of machines with certain functions; rather it is an expression of the social world"
- Nye (2006)
@@
Text-to-image (T2I) generative capabilities have advanced rapidly in recent years, exemplified by models such as Stable Diffusion-XL (Podell et al., 2023), Imagen 2 (Saharia et al., 2022), and DALLE-3 (Betker et al., 2023). As powerful tools for creative expression and communication, they have the potential to revolutionize numerous industries such as digital arts, advertising, and education. However, their widespread adoption across the globe raises important ethical and social considerations (Bird et al., 2023; Weidinger et al., 2023), in particular, in ensuring that these models work well for all people across the world (Qadri et al., 2023; Mim et al., 2024). While early T2I model evaluations focused on photo-realism (Saharia et al., 2022) and faithfulness/accuracy (Hu et al., 2023; Cho et al., 2024; Huang et al., 2023), recent work has demonstrated various societal biases that they reflect (Cho et al., 2023a; Bianchi et al., 2023; Luccioni et al., 2024). However, the predominantly mono-cultural development ecosystems of these models risks unequal representation of cultural awareness in them, potentially exacerbating existing technological inequalities (Prabhakaran et al., 2022). While the term “culture” has a myriad definitions across disciplines (Rapport & Overing, 2002), in this paper we focus on cultures formed within societies demarcated geographically through national boundaries (similar to Li et al. (2024c)), rather than cultures defined through organizational or other socio-demographic categories. This focus stems from our aim to assess global disparities in the capabilities of T2I models. Such disparities are shown to perpetuate harmful biases and stereotypes about cultures (Jha et al., 2024; Basu et al., 2023), as well as cause the erasure and suppression of sub- and co-cultures (Qadri et al., 2023), and limit their utility across geo-cultural contexts (Mim et al., 2024). While recent work built evaluation resources to detect various biases and stereotypes about different cultural groups that these models propagate (Jha et al., 2024; Basu et al., 2023), not much work has looked into how competent these models are in capturing the richness and diversity of various cultures.
@@
Gaps in such cultural competence may manifest primarily along two aspects of model generations: (i) cultural awareness: failure to recognize or generate the breadth of concepts/artifacts associated with a culture (Figure 1(c) and 1(d)), and (ii) cultural diversity: the tendency to adopt an oversimplified and homogenized view of a culture that associates (and generates) a narrow set of concepts/artifacts within that culture (Figure 1(b)) or across global cultures (Figure 1(a)). While the lack of cultural awareness in text to image models has been documented before (Hutchinson et al., 2022; Ventura et al., 2023), a major challenge in effectively assessing it at scale is the lack of resources that have a broad representation of cultural artifacts. Similarly, while dataset diversity has also been identified as an important part of the data-centric AI agenda (Oala et al., 2023) and has been investigated for text (Chung et al., 2023) and image modalities (Srinivasan et al., 2024; Dunlap et al., 2023), there has been limited focus on diversity of model generations (Lahoti et al., 2023), especially for T2I models. While works studying diversity of image generations focus on visual similarity (Hall et al., 2024; Zameshina et al., 2023), we study the diversity of generated cultural artifacts (aka cultural diversity).
@@
In this paper, we present CUBE: CUltural BEnchmark, a first-of-its-kind benchmark designed to facilitate the evaluation of cultural competence of text-to-image (T2I) models along two axes: cultural awareness and cultural diversity. We build this benchmark at the country level (in line with other recent work (Jha et al., 2024; Li et al., 2024c)), encompassing eight countries and representing three different concepts of cultural artifacts chosen as concepts of clear visual elements and hence of importance to T2I models. For each of these concepts, we employed a large-scale extraction strategy that leverages a Knowledge Graph (KG) augmented with a large language model (LLM) to build a broad-coverage compilation of country-specific concepts/artifacts to ground our evaluation. CUBE consists of a) CUBE-1K - a carefully curated subset of 1000 artifacts, made into prompts that enable evaluation of cultural awareness (Nguyen et al., 2023) via human annotation, and b) CUBE-CSpace - a collection of  300K cultural artifacts spanning the 8 countries and 3 concepts we consider. The methodology used to collect this resource is scalable to other concepts and geo-cultures. Furthermore, we introduce cultural diversity (CD) as a new evaluation component for T2I models, adapting the quality-weighted Vendi score (Nguyen & Dieng, 2024).
We detail the CUBE curation process in Section 3, cultural awareness evaluation in Section 4 and cultural diversity evaluation in Section 5.
To summarize, our main contributions are:
@@ 2 Related Work
In our discussion of related work, we focus on T2I evaluation metrics, culture in language technologies and vision, and T2I Benchmarks.
T2I Evaluation Metrics:
Culture in Language Technologies:
Culture in Vision:
T2I Benchmarks:
@@ 3 Construction of CUBE
Our benchmark aspires to enable reliable, trustworthy, and tangible measurement of text-to-image generative models for two distinct, yet complementary behaviors: cultural awareness (i.e., the model’s ability to reliably and accurately portray objects associated with a particular culture), and cultural diversity (i.e., the model’s ability to suppress oversimplified stereotypical depiction for an underspecified input that references a specific culture). One of the core prerequisites to meaningfully evaluate these aspects of cultural competence is a broad-coverage repository of cultural concepts/artifacts to ground such an evaluation. Inspired by previous work (Jha et al., 2024; Li et al., 2024c), we focus on geo-cultures (realized through the lens of national identity) to build such a repository, potentially extendable to other ways of categorizing culture, such as regions, religions, races, etc. We select eight countries from different geo-cultural regions across continents and the Global South-North divide: Brazil, France, India, Italy, Japan, Nigeria, Turkey, and USA. While we acknowledge that this list of countries is necessarily incomplete, and may result in a biased global sampling, future iterations of this work could include a wider range of countries for a more comprehensive evaluation.
Geo-culture Concept Cultural Artifacts
Japan Cuisine Ramen, Soba, Sushi, Katsu sandwich
India Art/Clothing Kurta, Lehanga Choli, Dhoti, Patola Saree
Table 2: Examples of cultural artifacts collected in CUBE-CSpace
@@
Finally, for each country-concept combination, we aim to construct grounding "concept spaces", which leads to a collection of 300K cultural artifacts, which we call CUBE-CSpace. This is an extensive compilation of concept space instances, also intended to be used as grounding for diversity evaluation. From this we create CUBE-1K: a much smaller, curated set of the 1000 artifacts across the 8 countries and 3 concepts - selected for relevance and popularity, intended to be used for testing cultural awareness. The country and concept wise split of CUBE-1K is presented in Table 8. In order to build CUBE, we adopt a Knowledge-Base (KB)-augmented LLM approach wherein we use graph-traversal on a pre-existing KB to extract a broad-coverage set of candidate cultural artifacts, followed by a self-critiquing LLM step to iteratively refine the repository.
@@
3.1 CUBE-CSpace
We use WikiData (Vrandečić & Krötzsch, 2014) as the KB to extract cultural artifacts, as it is the world’s largest publicly available knowledge base, with each entry intended to be supported by authoritative sources of information. We use the SLING framework to traverse the WikiData dump of April 2024222https://github.com/google/sling, by first manually identifying root nodes (see Table 9), a small seed set of manually selected nodes that represent the concept in question. For example, the node ’dish’ (WikiID: Q746549) is identified as a root node for the concept ’cuisine’. We then look for child nodes that lie along the ’instance of’ (P31) and ’subclass of’ (P279) edges; e.g. ’Biriyani’,(Q271555), a popular dish from India, is a child node of ’dish’ along the ’instance of’ edge. The child nodes that have the ’country-of-origin’ (P495) or the ’country’ (P17) are extracted as the cultural artifacts at the iteration. We recursively traverse the remaining nodes along the same edge classes in search of child nodes that satisfy these properties. For example ’bread’ (Q7802) is a child of ’dish’; since it is a generic food item, it doesn’t have the country (P17) or ’country-of-origin’ (P495) property. However, ’Filone’ (Q5449200) is a child of ’bread’ and has ’country-of-origin’ (P495) as Italy, which would be extracted at the step. The detailed extraction process is outlined in Algorithm 1. We extract artifacts from up to H=4 hop distances from the initial root nodes.
@@ 3.1 CUBE-CSpace
The above KB extraction process results in $\sim$500K collection of WikiData nodes, which is expected to have missing and inconsistent entries, owing to the noisy nature of WikiData (Kannen et al., 2023). We use GPT4-Turbo to filter out cultural artifacts that may not necessarily belong to a concept space. We take inspiration from existing self refinement (Madaan et al., 2023) and self critiquing (Lahoti et al., 2023) techniques to perform this step. Once we filter out the erroneous artifacts, we prompt GPT-4 to fill out popular missing artifacts from the cultural concept, similar to the diversity expansion application in (Lahoti et al., 2023). This filtering and completion process brings down the count to $\sim$300K entries, which forms the CUBE-CSpace. Table 2 presents some examples cultural artifacts extracted by this process.
@@ 3.2 CUBE-1K
As T2I models are primarily trained on English image-text pairs (Pouget et al., 2024), we expect them to struggle with visualizing long-tail artifacts from non-English-speaking cultures. To this end, CUBE-1K consists of prompts focusing on widely recognized artifacts, reflecting a model’s ability to capture mainstream cultural elements. The artifacts in CUBE-1K are a carefully curated subset of CUBE-CSpace. To ensure the inclusion of popular artifacts relevant to each country, we leverage the number of Google search results as a proxy for popularity. Specifically, we employ the Google Search API, utilizing the geolocation feature (’gl’ property) to tailor search results to a user located within the target country, thus capturing local popularity. We use this popularity estimate to sample artifacts for CUBE-1K. While search results serve as a useful proxy, we acknowledge they can be noisy, potentially inflated by the presence of popular keywords. Therefore, the final collection undergoes a manual verification process to ensure the quality and relevance of selected artifacts. CUBE-1K consists of 1000 prompts, spanning 8 countries and 3 concepts described above. Table 8 presents the distribution of artifacts across different countries in CUBE-1K. We use prompt templates designed to probe the models for cultural awareness, and specifically provide a negative prompt to obtain images of cultural artifacts with desirable qualities. Each prompt tests the T2I model’s ability to visualize a single cultural artifact. The prompt templates and negative prompt are provided in the Table 12.
@@ 4.1 Human Annotation
We recruited diverse groups of raters from each of the countries we consider. Each rater pool underwent comprehensive training and was also given a "golden set" of examples, as reference. Once training was complete, the raters proceeded to annotate the 1K prompts spanning the three concepts and the eight countries outlined in Table 3. Importantly, raters were instructed to focus solely on visual aspects when evaluating cultural relevance, while both image and text were considered for faithfulness. Realism was assessed independently of relevance and faithfulness. Detailed guidelines for each criterion can be found in Appendix B and the interface used in Figure 4.
We obtain high inter-rater agreement for the question on cultural relevance across all countries (all above 95%; see Table 11 in Appendix), suggesting that the question of whether an image is relevant to a particular culture is a relatively objective task. However, the question of faithfulness and realism yielded moderate to low agreement (especially for France and Turkey) among annotators (measured as Krippendorff’s $\alpha$, which is better suited for Likert scale ordinal values) in line with the relatively more complex and subjective nature of the task (see Table 13 for examples of edge cases).
@@ 4.2 Results
Table 3 shows the percentage of times our raters from each region deemed the images generated by each model to be culturally relevant (i.e., a yes answer to the first question in Annotation guidelines B). Both models yielded high scores for India, the USA, and France. Notably, Turkey and Nigeria received the lowest scores for Imagen, while Turkey and Brazil had the lowest scores for SDXL. All three of these countries belong to the Global South. Interestingly, these disparities are not uniform across models; for example, Brazil had a very high score for Imagen but the lowest score for SDXL. This suggests that the cultures marginalized by any particular model may depend on factors such as training data, reiterating the need for such cross-cultural benchmarks.
Model India Japan Italy USA Brazil France Turkey Nigeria
Imagen 98.4 90.8 86.8 95.0 94.3 95.2 68.1 73.6
SDXL 95.2 84.4 82.2 99.2 52.8 96.0 66.4 86.1
@@
Figure 3 shows some examples of faithfulness and realism scores for images that were deemed culturally relevant. In  3(a), the model was prompted to generate Pastel de angu from Brazilian cuisine and raters gave perfect score for both faithfulness and realism. In contrast, raters gave the lowest score of 1 for both aspects in  3(d), clearly identifying that it is neither faithful nor realistic. Similarly, the image of Sushi (3(b)) from Japanese cuisine got an faithfulness score of 5, but realism score of 1 with an observation: "The fish looks hard and made of glossy plastic.". Whereas, the image of lavallière from France is realistic but not faithful according to the raters. Table 4 presents the average consensus scores (and standard deviations) for both faithfulness and realism, as rated for each model across regions and concepts. Both Imagen 2 and SDXL exhibit substantial room for improvement in both faithfulness and realism. Consistent with prior observations (Table 3), both models achieve relatively lower scores for countries regarded as the Global South (such as Brazil, Turkey, and Nigeria), with this disparity particularly pronounced for faithfulness. On average, in comparison to faithfulness, realism scores are lower across geo-cultures. While Imagen generally outperforms SDXL, exceptions exist, such as art faithfulness in the USA where SDXL scores higher.

### v6

@@ 1 Introduction
Text-to-image (T2I) generative capabilities have advanced rapidly in recent years, exemplified by models such as Imagen 2 (Saharia et al., 2022), and DALLE-3 (Betker et al., 2023). As powerful tools for creative expression and communication, they have the potential to revolutionize numerous industries such as digital arts, advertising, and education. However, their widespread adoption across the globe raises important ethical and social considerations (Bird et al., 2023; Weidinger et al., 2023), in particular, in ensuring that these models work well for all people across the world (Qadri et al., 2023; Mim et al., 2024). While early T2I model evaluations focused on photo-realism (Saharia et al., 2022) and faithfulness(Hu et al., 2023; Cho et al., 2024; Huang et al., 2023), recent work has demonstrated various societal biases that they reflect (Cho et al., 2023a; Bianchi et al., 2023; Luccioni et al., 2024). However, the predominantly mono-cultural development ecosystems of these models risks unequal representation of cultural awareness in them, potentially exacerbating existing technological inequalities (Prabhakaran et al., 2022). While the term “culture” has a myriad definitions across disciplines (Rapport & Overing, 2002), in this paper we focus on cultures formed within societies demarcated geographically through national boundaries (similar to Li et al. (2024c)), rather than cultures defined through organizational or other socio-demographic categories. This focus stems from our aim to assess global disparities in the capabilities of T2I models. Such disparities are shown to perpetuate harmful stereotypes about cultures (Jha et al., 2024; Basu et al., 2023), as well as cause the erasure and suppression of sub- and co-cultures (Qadri et al., 2023), and limit their utility across geo-cultural contexts (Mim et al., 2024). While recent work has focussed on biases and stereotypes these models propagate (Jha et al., 2024; Basu et al., 2023), not much work has looked into how competent these models are in capturing the richness and diversity of various cultures.
@@
Gaps in cultural competence may manifest primarily along two aspects of model generations: (i) cultural awareness: failure to recognize or generate the breadth of concepts/artifacts associated with a culture (Figure 1(c) and 1(d)), and (ii) cultural diversity: the tendency to adopt an oversimplified and homogenized view of a culture that associates (and generates) a narrow set of concepts/artifacts within that culture (Figure 1(b)) or across global cultures (Figure 1(a)). While the lack of cultural awareness in text to image models has been documented before (Hutchinson et al., 2022; Ventura et al., 2023), a major challenge in effectively assessing it at scale is the lack of resources that have a broad representation of cultural artifacts. Similarly, while dataset diversity has also been identified as an important part of the data-centric AI agenda (Oala et al., 2023) and has been investigated for text (Chung et al., 2023) and image modalities (Srinivasan et al., 2024; Dunlap et al., 2023), there has been limited focus on diversity of model generations (Lahoti et al., 2023), especially for T2I models. While works studying diversity of image generations focus on visual similarity (Hall et al., 2024; Zameshina et al., 2023), we study the diversity of generated cultural artifacts (aka cultural diversity).
@@
In this paper, we present CUBE: CUltural BEnchmark, a first-of-its-kind benchmark designed to facilitate the evaluation of cultural competence of T2I models along two axes: cultural awareness and cultural diversity. We build this benchmark at the country level (in line with recent works (Jha et al., 2024; Li et al., 2024c)), encompassing eight countries and representing three different concepts of cultural artifacts chosen as concepts of clear visual elements, and hence of importance to T2I models. We employed a large-scale extraction strategy that leverages a Knowledge Graph (KG) augmented with a Large Language Model (LLM) to build a broad-coverage compilation of country-specific artifacts to ground our evaluation. CUBE consists of a) CUBE-1K - a carefully curated subset of 1000 artifacts, made into prompts that enable evaluation of cultural awareness (Nguyen et al., 2023), and b) CUBE-CSpace - a collection of $\sim$300K cultural artifacts spanning the 8 countries and 3 concepts we consider, with a potential to be scaled to other concepts and countries. Furthermore, we introduce cultural diversity (CD) as a new evaluation component for T2I models, adapting the quality-weighted Vendi score (Nguyen & Dieng, 2024).We detail the CUBE curation process in Section 3, cultural awareness evaluation in Section 4 and cultural diversity evaluation in Section 5. To summarize, our main contributions are:
@@ 2 Related Work
In our discussion of related work, we focus on T2I evaluation and culture in large models.
T2I Evaluation.
Inception Score (Salimans et al., 2016) and Frechet Inception Distance (Heusel et al., 2018) focus on similarity of generated images to real ones, also called realism. Metrics like DSG (Cho et al., 2024) and VQAScore (Lin et al., 2024) measure the prompt-image alignment, also called faithfulness. Other metrics such as ImageReward (Xu et al., 2023), PickScore (Kirstain et al., 2023), and HPSv2 (Wu et al., 2023) fine-tune vision-language models on human ratings to better align with human preferences. There have been recent works on bias and fairness evaluation (Feng et al., 2022; Naik & Nushi, 2023; Zhang et al., 2023; Jha et al., 2024) of T2I models. There have also been efforts to build comprehensive evaluation benchmarks aimed at tracking the progress of model capabilities over time, focusing on tasks such as realism, text faithfulness, and compositional abilities. These benchmarks, such as DrawBench (Saharia et al., 2022), CC500 (Feng et al., 2023), T2I-CompBench (Huang et al., 2023), TIFA v1.0 (Hu et al., 2023), DSG-1k (Cho et al., 2024), GenEval (Ghosh et al., 2023), and GenAIBench (Lin et al., 2024) employ diverse prompts and metrics to assess factors such as image-text coherence, perceptual quality, attribute binding, faithfulness, semantic competence, and compositionality, to list a few.
Culture in Language Technologies.
Culture in Vision.
@@ 3 Construction of CUBE
Our benchmark aspires to enable reliable, trustworthy, and tangible measurement of text-to-image generative models for two distinct yet complementary behaviors: cultural awareness (i.e., the model’s ability to reliably and accurately portray objects associated with a particular culture), and cultural diversity (i.e., the model’s ability to suppress oversimplified stereotypical depiction for an underspecified input that references a specific culture). One of the core prerequisites to meaningfully evaluate these aspects of cultural competence is a broad-coverage repository of cultural artifacts to ground such an evaluation. Inspired by previous work (Jha et al., 2024; Li et al., 2024c), we focus on geo-cultures (realized through the lens of national identity) to build such a repository, potentially extendable to other ways of categorizing culture, such as regions, religions, races, etc. We select eight countries from different geo-cultural regions across continents and the Global South-North divide: Brazil, France, India, Italy, Japan, Nigeria, Turkey, and USA. While we acknowledge that this list of countries is necessarily incomplete, and may result in a biased global sampling, future iterations of this work could include a wider range of countries for a more comprehensive evaluation.
@@
Finally, for each country-concept combination, we aim to construct grounding "concept spaces", which leads to a collection of $\sim$300K cultural artifacts, which we call CUBE-CSpace. This is an extensive compilation of concept space instances, also intended to be used as grounding for diversity evaluation. From this, we create CUBE-1K: a much smaller, curated set of the 1000 artifacts across the 8 countries and 3 concepts - selected for relevance and popularity, intended to be used for testing cultural awareness. The country and concept wise split of CUBE-1K is presented in Table 7. In order to build CUBE, we adopt a Knowledge-Base (KB)-augmented LLM approach wherein we use graph-traversal on a pre-existing KB to extract a broad-coverage set of candidate cultural artifacts, followed by a self-critiquing LLM step to iteratively refine the repository.
3.1 CUBE-CSpace
@@
We use WikiData (Vrandečić & Krötzsch, 2014) as the KB to extract cultural artifacts, as it is the world’s largest publicly available knowledge base, with each entry intended to be supported by authoritative sources of information. We use the SLING framework [^2] https://github.com/google/sling to traverse the WikiData dump of April 2024, by first manually identifying root nodes (see Table 8), a small seed set of manually selected nodes that represent the concept in question. For example, the node ’dish’ (WikiID: Q746549) is identified as a root node for the concept ’cuisine’. We then look for child nodes that lie along the ’instance of’ (P31) and ’subclass of’ (P279) edges; e.g. ’Biriyani’,(Q271555), a popular dish from India, is a child node of ’dish’ along the ’instance of’ edge. The child nodes that have the ’country-of-origin’ (P495) or the ’country’ (P17) are extracted at the iteration. We recursively traverse the remaining nodes along the same edge classes in search of child nodes that satisfy these properties. For example ’bread’ (Q7802) is a child of ’dish’; since it is a generic food item, it doesn’t have the ’country-of-origin’ (P495) property. However, ’Filone’ (Q5449200) is a child of ’bread’ and has ’country-of-origin’ (P495) as Italy, which would be extracted at the step. We outline the extraction process in Algorithm 1. In practise, we iterated for H=4 hops and have detailed considerations in Appendix C.4.
@@
The above KB extraction process results in $\sim$500K collection of WikiData nodes, which is expected to have missing and inconsistent entries, owing to the noisy nature of WikiData (Kannen et al., 2023). We use GPT4-Turbo to filter out cultural artifacts that may not necessarily belong to a concept space, taking inspiration from existing self refinement (Madaan et al., 2023) and self critiquing (Lahoti et al., 2023) techniques. Once we filter out the erroneous artifacts, we prompt GPT-4 to fill out popular missing artifacts from the cultural concept, similar to the diversity expansion application in (Lahoti et al., 2023). This filtering and completion process brings down the count to $\sim$300K entries, which forms the CUBE-CSpace. Table 2 presents some examples cultural artifacts extracted by this process.
@@ 3.2 CUBE-1K
As T2I models are primarily trained on English image-text pairs (Pouget et al., 2024), we expect them to struggle with visualizing artifacts from non-English-speaking cultures. To this end, CUBE-1K consists of prompts focusing on widely recognized artifacts, reflecting a model’s ability to capture mainstream cultural elements. The artifacts in CUBE-1K are a carefully curated subset of CUBE-CSpace. To ensure the inclusion of popular artifacts relevant to each country, we leverage the number of Google search results as a proxy for popularity. Specifically, we employ the Google Search API, utilizing the geolocation feature (’gl’ property) to tailor search results to a user located within the target country, thus capturing local popularity. We use this popularity estimate to sample artifacts for CUBE-1K. While search results serve as a useful proxy, we acknowledge they can be noisy, potentially inflated by the presence of popular keywords. Therefore, the final collection undergoes a manual verification process (detailed in Appendix C.4) to ensure relevance of selected artifacts. CUBE-1K consists of 1000 prompts, spanning 8 countries and 3 concepts described above. Table 7 presents the distribution of artifacts across different countries in CUBE-1K. We use prompt templates designed to probe the models for cultural awareness, along with a negative prompt (Hao et al., 2023) to obtain images with desired qualities. Each prompt tests the model’s ability to visualize a single artifact. The prompt templates and negative prompt are provided in the Table 13 (in Appendix).
@@ 4 Evaluating Cultural Awareness
Geo-culture Concept Cultural Artifacts
Japan Cuisine Ramen, Soba, Sushi, Katsu sandwich
India Art/Clothing Kurta, Lehanga Choli, Dhoti, Patola Saree
Table 2: Examples of cultural artifacts collected in CUBE-CSpace
@@ 4.1 Human Annotation
We recruited diverse groups of raters from each of the countries we consider. Each rater pool underwent comprehensive training and was also given a "golden set" of examples, as reference. Once training was complete, the raters proceeded to annotate the 1K prompts spanning the three concepts and the eight countries outlined in Table 3. Raters were instructed to focus on both the image and text when evaluating cultural relevance, and solely the image for realism. Detailed guidelines for each criterion (Appendix D), the inter-annotator agreement (Appendix E), and the interface used for human annotation (Figure 4) can be found in Appendix.
@@ 4.2 Results
Figure 3 presents examples of faithfulness and realism scores for images that were deemed culturally relevant. In  3(a), the model was prompted to generate Pastel de angu from Brazilian cuisine and raters gave perfect score for both faithfulness and realism. In contrast, raters gave the lowest score of 1 for both aspects in  3(d), clearly identifying that it is neither faithful nor realistic. Similarly, the image of Sushi (3(b)) from Japanese cuisine got an faithfulness score of 5, but realism score of 1 with an observation: "The fish looks hard and made of glossy plastic.". Whereas, the image of lavallière from France is realistic but not faithful, according to the raters.
@@
Table 3 presents the average consensus scores (and standard deviations) for both faithfulness and realism, as rated for each model across regions and concepts. Both Imagen 2 and SDXL exhibit substantial room for improvement in both faithfulness and realism. Both models achieve relatively lower scores for countries regarded as the Global South (such as Brazil, Turkey, and Nigeria), with this disparity particularly pronounced for faithfulness. On average, in comparison to faithfulness, realism scores are lower across geo-cultures. While Imagen generally outperforms SDXL, exceptions exist, such as art faithfulness in the USA where SDXL scores higher. Table 12 (in Appendix) shows the percentage of times our raters from each region deemed the images generated by each model to be culturally relevant (i.e., a yes answer to the first question in Annotation guidelines D) showing non-uniform disparities across models and cultures. This suggests that the cultures marginalized by any particular model may depend on factors such as training data, reiterating the need for such cross-cultural benchmarks.